# Measurement-induced physics and stochastic dynamics in systems of single-body quasiparticle and many-body microscopic cases

### v2

@@ V Stochastic evolution with Markovian and non-Markovian dynamics
While the nonlocal correlations corresponds to the nonlinear one, like the von Neumann entropy and von Neumann correlation. Only the latter is free from Lieb-Robinson bound and deeply affected by the local quantum channel, e.g., the ensemble average over a large number of quantum jumps such that the local quantum channel has a large enough width to take effect while a narrow quantum channel (mixed state) is in the sense that it can be purified to a symmetric local unitary operator[78], and local correlations are avoided by the statistical mixture which generally corresponds to weak symmetry[39, 77, 78]. Also, this is also in agree with the suppression on area-law entanglement by the quantum jumps which plays the role of perturbation.
Further, in the case neither dominated by the classical dynamics or reaching the thermalization, the quantum information over large distance is allowed in inhomogeneous steady state with extensive long-range entanglement or the postquenched homogeneous integrable state[74].
VI Conclusion
Generally speaking, the steady state of $mm$-case (consistent with the open boundary condition) is only due to the skin effect, where the many-body state (without exponential localization) holds global fluctuation due to the Hermiticity and the long-time behavior can be well predicted by diagonalization. This is in contrast to the localized state in $sq$-case (consistent with the periodic boundary condition) where the single-particle eigenstates exponentially decay with the localization length independent of system size[56, 94, 10, 124, 125, 130], and thus the long-time behavior is harder to predict by approximately diagonalization due to the non-Hermiticity. Zeno effect and confined light cone[131, 96] are unique to the $sq$ case, while Hilbert space fragmentation[154, 121] and linearized bosonization are unique to the $mm$ case, and the topological islands[125] is unique to the intermedia regime with finite strength of disorder but not yet reaches the ensemble-averaged result.
VII Appendix: Method
We define the densities of the two classes as $\rho_{mm}:=\sum_{i}|\psi_{i}\rangle\langle\psi_{i}|$, $\rho_{sq}:=\sum_{i}|\psi_{i}\rangle\langle\psi^{*}_{i}|$. The reduced form can be seem by the state of the combined system environment $|\rho_{SE}\rangle=\sum_{i}|\psi_{i}\rangle|\mathcal{E}_{i}\rangle$ with $\langle\mathcal{E}_{i}|\mathcal{E}_{j}\rangle=\langle\mathcal{E}_{i}|\mathcal{E}_{j}^{*}\rangle=\delta_{ij}$. The complex eigenstates are prepared aimed to obtaining the non-Hermitian operator ${\bf\Psi}$ whose elements follow the distribution of Gaussian ensemble, upto second and fourth moments, and reads, when consider two real eigenvalues, ${\bf\Psi}=\lambda_{\alpha}{\bf\Psi}_{\alpha}+\lambda_{\beta}{\bf\Psi}_{\beta}$ For a general route of construction, see Refs.[150]. The densities $D_{mm/sq}(t)=e^{i\rho_{mm/sq}t}(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}|)$, whose square of trace is the Loschmidt echo, is important in revealing the difference between $sq$ and $mm$ cases. The initial state for both cases are pure state and unentangled. Available convex sum form for densities in $mm$-case (diagonal in the Schmidt basis[156]) agrees with the absence of topological order therein. Environmental decoherence effect (off-diagonal fluctuation) emerges when $\langle\mathcal{E}_{i}|\mathcal{E}_{j}\rangle\neq 0$ for $i\neq j$, and this will suppress the entanglement and cause the gapless structure for $mm$-case and early stage of $sq$-case. Such decoherence and off-diagonal fluctuation will be suppressed by larger system size in $sq$-case, and decay with time toward the gapped (area law). Also, the entropy for product state $(\sum_{i}|\psi_{i}\rangle)(\sum_{i}\langle\psi_{i}^{*}|)$ converges to $1/2^{N}\approx 0.0039$ where $N=8$ here is the maximal number of independent states. While such decoherence as well as stochastic dynamics is persistent in $mm$-case, due to the restricted Hilbert space and disorders which guarantee the accessibility of (maximally mixed) infinite-temperature featureless equilibrium state in long-time limit.
As we shown in Fig.2(a), for $sq$-case, the entanglement entropy begin with a linear increase following the volume law, and then power law growth to maximal value. The enhancement of entanglement in this stage mainly originates from the increase of correlation range, during which process the non-Gaussian interactions be suppressed due to the global dephasing. After the maximal entropy, there is firstly a obvious power law decay (critical phase) and them a unobtrusive logarithmical increase, but the whole decay process toward the long-time (nonequilibrium) steady state can be enveloped by a exponentially fluctuating decay, as agreed by the emergent gapped mode and single-particle character. The entropy starts to decay when the entanglement production due to the enlongated range of interactions cannot catch up with the annihilation due to the weakened interaction strength. The latter stage of $sq$-case has same similarities with the dynamical mean field theory, i.e., the suppressed entropy and fluctuations. But for DMFT, instead of the global symmetry, it is the local Markovian drive which causes the fast dephasing, and the non-Markovian bath algebraically decay with the increasing lattice connectivity (in a next-nearest interacting model)[152, 148]. Our result also shows that, larger system size (integer multiples of a complete basis to ensure the global dephsing) cause faster relaxation toward the nonequilibrium steady state at finite temperature, and also faster growth to the maximal entropy. This is because larger system size induces stronger suppression on the non-Gaussian interaction (and the associate averaging effect), which is the same with the effect of denser measurement[136]. The exponential decay in latter stage reflecting the area law region, is faster for smaller system size and thus larger gap of Liouville superoperator, in agree with Ref.[151].
[…]

### v4

@@ VII Appendix: Stochastic evolution with Markovian and non-Markovian dynamics
While the nonlocal correlations corresponds to the nonlinear one, like the von Neumann entropy and von Neumann correlation. Only the latter is free from Lieb-Robinson bound and deeply affected by the local quantum channel, e.g., the ensemble average over a large number of quantum jumps such that the local quantum channel has a large enough width to take effect while a narrow quantum channel (mixed state) is in the sense that it can be purified to a symmetric local unitary operator[78], and local correlations are avoided by the statistical mixture which generally corresponds to weak symmetry[39, 77, 78]. Also, this is also in agree with the suppression on area-law entanglement by the quantum jumps which plays the role of perturbation. Further, in the case neither dominated by the classical dynamics or reaching the thermalization, the quantum information over large distance is allowed in inhomogeneous steady state with extensive long-range entanglement or the postquenched homogeneous integrable state[74].
VIII Appendix.B: Effect from Lindblad dynamics to $sq$-case
For non-Hermitian and nonunitary observables $O_{nH}=e^{-i\rho_{sq}t}$ ($t\neq 0$), $|\langle\rho_{0}|O_{nH}|\rho_{0}\rangle|=|\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle|\approx|\langle\rho^{\prime}_{0}|O_{nH}|\rho^{\prime}_{0}\rangle|$, and $\langle\rho^{\prime}_{0}|R_{sq}|\rho_{0}\rangle/\langle\rho^{\prime}_{0}|\rho_{0}\rangle\neq 1$, which means the effect of measurement apparatus keep finite but small during the evolution (also implies the perturbative treatment is invalid here). Then the effect of imaginary part related to the measurement-induced interaction can be estimated from the difference $\langle\rho_{0}|O_{nH}|\rho_{0}\rangle-\langle\rho_{0}|UO_{nH}U^{{\dagger}}|\rho_{0}\rangle$ where $U$ is nonunitary such that $[O_{nH},{\rm ln}U]\neq 0$ and ${\rm Tr}[UO_{nH}U^{{\dagger}}-O_{nH}]=0$.
Figure 8: Von Neumann entropy $S=-{\rm ln}{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]$ for different system sizes.
We show the von Neumann entropy $S=-{\rm ln}{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]$ in Fig.8. Compares to the von Neumann entropy shwon in Fig.1(a), we see that the the peak at intermediate time vanishes and the entropy become monotonic and smooth over the whole evolution.
Next we consider the effect from weak interaction and the one body loss $\gamma$, which is also the decoherence function, the dynamics of time-dependent density $R_{sq}$ can be described by the Lindblad function especially in long-time limit,
$\frac{dR_{sq}}{dt}=-i(H_{int}R_{sq}-R^{{\dagger}}_{sq}H^{{\dagger}}_{int})+\frac{\gamma}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)+\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right),$ (5)
where ${\bf I}$ is the $N\times N$ identity matrix, and the interacting Hamiltonian reads $H_{int}=g|\rho_{0}\rangle\langle\rho^{\prime}_{0}|+h.c.$. Then we numerically found that, the dynamics of expectation of $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle)$ can be well described by
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\dot{R}_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle)$ (6)
$=$ $-\frac{ig{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}-\frac{ig{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$
$-\frac{ig{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}-\frac{ig{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$
$+\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$
$+\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}+\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{\gamma}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}.$
We have the following relations at arbitary $t$,
$\frac{-i{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=-i{\rm Tr}[(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}],$ (7)
$\frac{-i{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=-i{\rm Tr}[(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}],$
$\frac{-i{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=-i{\rm Tr}[(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}],$
$\frac{-i{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|(|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=-i{\rm Tr}[(|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}-R^{{\dagger}}_{sq}|\rho_{0}\rangle\langle\rho^{\prime}_{0}|)]=2\ {\rm Im}\ {\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}].$
Note that $\ {\rm Im}\ {\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]=-\ {\rm Im}\ {\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]$.
At long-time where $U_{sq}$ is trace preserving and $1/{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|=1/{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|=1/N$, there is symmetry for the exchange between states before and after measurement,
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]({\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]-1)}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]({\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]-1)}=-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=-\frac{1}{N},$ (8)
$\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{2{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{1}{2N},$
Note the following relations,
$\frac{{\rm Tr}[R_{sq}]}{N}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{N^{2}}$ (9)
$=-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]({\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]-1)}=-\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]({\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]-1)}\sim e^{-t},$
$||\ |\rho_{0}\rangle\langle\rho_{0}|\ ||_{2}=||\ |\rho^{\prime}_{0}\rangle\langle\rho_{0}|\ ||_{2}=||\ |\rho_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}=||\ |\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\ ||_{2}=\frac{1}{N}||\ |\rho_{0}\rangle\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle\langle\rho_{0}|\ ||_{2}$
$={\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]={\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]=N\ (\forall\ t),$
as shown in Fig.9(a) where $\langle R_{sq}\rangle_{1}$, $\langle R_{sq}\rangle_{2}$, $\langle\mathcal{D}\rangle_{1}$ represent the second, third, and fourth term of the first expression of Eq.9. While $\langle R_{sq}\rangle_{3}:=\frac{1}{N}|{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]|$ and $\langle R_{sq}\rangle_{4}:=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]}$ are shown inFig.9(a)-(b). At long-time where $U_{sq}$ still be trace preserving,
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{||[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]||_{2}}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{||[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]||_{2}}=\frac{1}{N},$ (10)
where the 2-norm and trace norm satisfy $||[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]||_{2}=2||[|\rho_{0}\rangle\langle\rho^{\prime}_{0}|]||_{tr}=N$.
Figure 9: (a) Blue, orange and green dots represent the expectations $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}$, $\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}$ and $\bigg{|}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{||[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]||_{2}}\bigg{|}$, respectively. (b)Blue, orange and green dots represent the real part, imaginary part, and absolute value of $\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho_{0}|]}$.
While at long-time limit $t\rightarrow\infty$ where $U_{sq}$ is nomore trace preserving and $1/{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|=1/N,1/{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|=0$, we have
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]({\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]-1)}=-\frac{1}{N},$ (11)
$\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]}=\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\frac{1}{N},$
$\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]({\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|]-1)}=0,$
$\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\frac{1}{2}\left(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\}\right)]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}=\infty,$
To obtain a long-time-stable result, we restrict the dissipative part of the dynamics of $R_{sq}$ (Eq.(5)) to only containing the state before nonunitary evolution (measurement), i.e., $|\rho_{0}\rangle$. then we have
$\lim_{t\rightarrow\infty}\left[\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\dot{R}_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}+(|\rho_{0}\rangle\leftrightarrow|\rho^{\prime}_{0}\rangle)\right]=\gamma\frac{2-N}{N},$ (12)
which is nearly $-\gamma$ in the large-$N$ limit. The long-time limit of $\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\dot{R}_{sq}]}{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|]}$ turns to be the same result. This result is consistent with Refs., such that the dissipative part of the dynamics will not generates time-dependence to the postselection probability at long-time limit, as long as the decoherence function is time-independent.
Next we consider the effect from Lindblad dynamics to the time evolution of von Neumann entropy $S=-{\rm ln}{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]$,
$\dot{S}$ $=\frac{-1}{{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]}{\rm Tr}[\frac{dR_{sq}}{dt}|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|+R_{sq}|\rho_{0}\rangle\langle\rho_{0}|\frac{dR_{sq}}{dt}|\rho_{0}\rangle\langle\rho_{0}|]$ (13)
$=\frac{-2}{{\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}]}{\rm Tr}[\frac{dR_{sq}}{dt}|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|].$
We found that the unitary part in the Lindblad function shows to be irrelavant to introduce large randomness determined by the interaction strength $g$, thus here we consider only the role of dissipative part, i.e., $g\ll\gamma$. Then we define
$f_{1}:={\rm Tr}[\frac{1}{2}(2\frac{\langle\rho_{0}|R_{sq}|\rho_{0}\rangle}{\langle\rho_{0}|\rho_{0}\rangle}{\bf I}-\{|\rho_{0}\rangle\langle\rho_{0}|,R_{sq}\})|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|],$ (14)
$f_{2}:={\rm Tr}[\frac{1}{2}(2\frac{\langle\rho^{\prime}_{0}|R_{sq}|\rho^{\prime}_{0}\rangle}{\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle}{\bf I}-\{|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|,R_{sq}\})|\rho_{0}\rangle\langle\rho_{0}|R_{sq}|\rho_{0}\rangle\langle\rho_{0}|],$
$f_{3}:={\rm Tr}[(R_{sq}|\rho_{0}\rangle\langle\rho_{0}|)^{2}].$
As shown in Fig.10, $f_{1}$ exhibits the same exponential decay with $f_{3}$, while $f_{2}$ decay with fluctuations, and $f_{2}$ is bound by $f_{1}$ from below. Further, we found that, by renormalizing the functions such that their initial value equals 1, we have $\tilde{f_{1}}=\tilde{f_{3}}$. Thus the $f_{1}$ does not contributions to the time-dependence of the entropy, while $f_{2}$ soly leads to fluctuations with decaying amplitude. Roughly speaking, $\langle\rho^{\prime}_{0}|\rho_{0}\rangle$ contributes to the fluctuation last for a long time (as long as $U_{sq}$ be trace preserving), while $\langle\rho^{\prime}_{0}|\rho^{\prime}_{0}\rangle$ and $\langle\rho_{0}|\rho_{0}\rangle$ contribute to the exponential decay with time. To understand this, we consider the correlation function between the two states (before and after measurement)
$\mathcal{C}_{1}(t):={\rm Re}\left(\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}-\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}\frac{{\rm Tr}[|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}\right)={\rm Re}\frac{{\rm Tr}[|\rho_{0}\rangle\langle\rho_{0}|\rho^{\prime}_{0}\rangle\langle\rho^{\prime}_{0}|R_{sq}]}{{\rm Tr}[R_{sq}]}-1,$ (15)
[…]